# Chain-linked Multiple Matrix Integration via Embedding Alignment

### v2

@@ Appendix D Technical Lemmas
Eq. (D.15) and Eq. (D.16) implies $\|\mathbf{A}^{[h]}-\mathbf{P}\|\lesssim q_{i}^{-1/2}(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})n_{i}^{1/2}$ and $\lambda_{d}(\mathbf{A}^{[h]})\asymp\lambda_{i,\min}$ with high probability. Then by Wedin’s $\sin\Theta$ Theorem (see e.g., Theorem 4.4 in Stewart and Sun (1990)) we have
$\|\sin\Theta(\hat{\mathbf{U}}^{[h]},\mathbf{U})\|\leq\frac{\|\mathbf{A}^{[h]}-\mathbf{P}\|}{\lambda_{d}(\mathbf{A}^{[h]})-\lambda_{d+1}(\mathbf{P})}\lesssim\frac{(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})n_{i}^{1/2}}{q_{i}^{1/2}\lambda_{i,\min}}$
with high probability. Let $\mathbf{W}^{[h]}$ be the orthogonal Procrustes alignment between $\hat{\mathbf{U}}^{[h]}$ and $\mathbf{U}$. Then
$\|\hat{\mathbf{U}}^{[h]\top}\mathbf{U}-\mathbf{W}^{[h]}\|\leq\|\sin\Theta(\hat{\mathbf{U}}^{[h]},\mathbf{U})\|^{2}\lesssim\frac{(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})^{2}n_{i}}{q_{i}\lambda_{i,\min}^{2}},$ (D.22)
@@
Now let $\mathbf{Z}^{[h]}=\hat{\mathbf{U}}^{[h]}\hat{\mathbf{U}}^{[h]\top}\mathbf{U}-\mathbf{U}$. By Eq. (D.22) and Lemma D.5 we obtain
$\|\mathbf{Z}^{[h]}\|_{2\to\infty}$ $\leq\|\hat{\mathbf{U}}^{[h]}\hat{\mathbf{U}}^{[h]\top}\mathbf{U}-\hat{\mathbf{U}}^{[h]}\mathbf{W}^{[h]}\|_{2\to\infty}+\|\hat{\mathbf{U}}^{[h]}\mathbf{W}^{[h]}-\mathbf{U}\|_{2\to\infty}$
$\leq\|\hat{\mathbf{U}}^{[h]}\|_{2\to\infty}\cdot\|\hat{\mathbf{U}}^{[h]\top}\mathbf{U}-\mathbf{W}^{[h]}\|+\|\hat{\mathbf{U}}^{[h]}\mathbf{W}^{[h]}-\mathbf{U}\|_{2\to\infty}$
@@
with high probability, where the final inequality follows from the fact that, under the conditions in Eq. (C.3) we have $\frac{(\|\mathbf{P}\|_{\max}+\sigma_{i})n_{i}^{1/2}}{q^{1/2}\lambda_{i,\min}\log^{1/2}n_{i}}\lesssim 1$.
Finally, as $\mathbf{e}_{h}$ and $\mathbf{Z}^{[h]}$ are independent, by Lemma D.6 and the assumption $q_{i}n_{i}\gtrsim\log^{2}n_{i}$ we have
@@
Lemma D.9.
Consider the setting in Lemma D.4, we then have
$\|\mathbf{E}^{(i)}(\hat{\mathbf{U}}^{(i)}\mathbf{W}_{\mathbf{U}}^{(i)}-\mathbf{U}^{(i)})\|_{2\to\infty}\lesssim\frac{d^{1/2}(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})^{2}n_{i}^{1/2}\log n_{i}}{q_{i}\lambda_{i,\min}}$
@@
$\mathbf{e}_{h}^{\top}(\hat{\mathbf{U}}\mathbf{W}_{\mathbf{U}}-\mathbf{U})$ $=\mathbf{e}_{h}^{\top}(\hat{\mathbf{U}}\mathbf{W}_{\mathbf{U}}-\mathbf{U})\mathbf{W}_{\mathbf{U}}^{\top}(\mathbf{W}_{\mathbf{U}}-\hat{\mathbf{U}}^{\top}\mathbf{U})+\mathbf{e}_{h}^{\top}\mathbf{U}\mathbf{W}_{\mathbf{U}}^{\top}(\mathbf{W}_{\mathbf{U}}-\hat{\mathbf{U}}^{\top}\mathbf{U})$ (D.23)
$+\mathbf{e}_{h}^{\top}(\hat{\mathbf{U}}\hat{\mathbf{U}}^{\top}-\hat{\mathbf{U}}^{[h]}\hat{\mathbf{U}}^{[h]\top})\mathbf{U}+\mathbf{e}_{h}^{\top}(\hat{\mathbf{U}}^{[h]}\hat{\mathbf{U}}^{[h]\top}\mathbf{U}-\mathbf{U}).$
We now bound all terms on the right hand side of Eq. (D.23). For $\mathbf{e}_{h}^{\top}(\hat{\mathbf{U}}\mathbf{W}_{\mathbf{U}}-\mathbf{U})\mathbf{W}_{\mathbf{U}}^{\top}(\mathbf{W}_{\mathbf{U}}-\hat{\mathbf{U}}^{\top}\mathbf{U})$, by Lemma D.2 we have
$\|\mathbf{e}_{h}^{\top}(\hat{\mathbf{U}}\mathbf{W}_{\mathbf{U}}-\mathbf{U})\mathbf{W}_{\mathbf{U}}^{\top}(\mathbf{W}_{\mathbf{U}}-\hat{\mathbf{U}}^{\top}\mathbf{U})\|$ $\lesssim\|\mathbf{e}_{h}^{\top}(\hat{\mathbf{U}}\mathbf{W}_{\mathbf{U}}-\mathbf{U})\|\cdot\|\mathbf{W}_{\mathbf{U}}-\hat{\mathbf{U}}^{\top}\mathbf{U}\|$ (D.24)
$\lesssim\|\mathbf{e}_{h}^{\top}(\hat{\mathbf{U}}\mathbf{W}_{\mathbf{U}}-\mathbf{U})\|\cdot\frac{(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})^{2}n_{i}}{q_{i}\lambda_{i,\min}^{2}}$
$=O(\|\mathbf{e}_{h}^{\top}(\hat{\mathbf{U}}\mathbf{W}_{\mathbf{U}}-\mathbf{U})\|)$
with high probability as $\frac{(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})n_{i}^{1/2}}{q_{i}^{1/2}\lambda_{i,\min}}\lesssim 1$ as implied by Eq. (C.3).
For $\mathbf{e}_{h}^{\top}\mathbf{U}\mathbf{W}_{\mathbf{U}}^{\top}(\mathbf{W}_{\mathbf{U}}-\hat{\mathbf{U}}^{\top}\mathbf{U})$, by Lemma D.1 and Lemma D.2 we have
$\|\mathbf{e}_{h}^{\top}\mathbf{U}\mathbf{W}_{\mathbf{U}}^{\top}(\mathbf{W}_{\mathbf{U}}-\hat{\mathbf{U}}^{\top}\mathbf{U})\|$ $\leq\|\mathbf{E}\mathbf{U}\|_{2\to\infty}\cdot\|\mathbf{W}_{\mathbf{U}}-\hat{\mathbf{U}}^{\top}\mathbf{U}\|$ (D.25)
@@
For $\mathbf{e}_{h}^{\top}(\hat{\mathbf{U}}\hat{\mathbf{U}}^{\top}-\hat{\mathbf{U}}^{[h]}\hat{\mathbf{U}}^{[h]\top})\mathbf{U}$, by Lemma D.1 and Lemma D.7 we have
$\|\mathbf{e}_{h}^{\top}(\hat{\mathbf{U}}\hat{\mathbf{U}}^{\top}-\hat{\mathbf{U}}^{[h]}\hat{\mathbf{U}}^{[h]\top})\mathbf{U}\|$ $\leq\|\mathbf{E}\|\cdot 2\|\sin\Theta(\hat{\mathbf{U}}^{[h]},\hat{\mathbf{U}})\|$ (D.26)
@@
For $\mathbf{e}_{h}^{\top}(\hat{\mathbf{U}}^{[h]}\hat{\mathbf{U}}^{[h]\top}\mathbf{U}-\mathbf{U})$, by Lemma D.8 we have
$\|\mathbf{e}_{h}^{\top}(\hat{\mathbf{U}}^{[h]}\hat{\mathbf{U}}^{[h]\top}\mathbf{U}-\mathbf{U})\|\lesssim\frac{d^{1/2}(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})^{2}n_{i}^{1/2}\log n_{i}}{q_{i}\lambda_{i,\min}}$ (D.27)
@@
Combining Eq. (D.23), Eq. (D.24), …, Eq. (D.27) we finally obtain
$\|\mathbf{e}_{h}^{\top}(\hat{\mathbf{U}}\mathbf{W}_{\mathbf{U}}-\mathbf{U})\|$ $\lesssim\frac{d^{1/2}(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})^{2}n_{i}^{1/2}\log n_{i}}{q_{i}\lambda_{i,\min}}$
with high probability as $\frac{(\|\mathbf{P}\|_{\max}+\sigma_{i})n_{i}^{1/2}}{q_{i}^{1/2}\lambda_{i,\min}\log^{1/2}n_{i}}\lesssim 1$ as implied by Eq. (C.3). ∎
D.5 Proof of Lemma C.2
@@
$\mathbf{W}^{(i)\top}\mathbf{W}^{(i,j)}\mathbf{W}^{(j)}=\underset{\mathbf{O}\in\mathbb{O}_{d}}{\arg\min}\|\hat{\mathbf{X}}^{(i)}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}\mathbf{W}^{(i)}\mathbf{O}-\hat{\mathbf{X}}^{(j)}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}\mathbf{W}^{(j)}\|_{F}.$
@@
$\mathbf{F}:=\mathbf{W}^{(i)\top}(\hat{\mathbf{X}}^{(i)}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle})^{\top}\hat{\mathbf{X}}^{(j)}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}\mathbf{W}^{(j)}-\mathbf{X}_{\mathcal{U}_{i}\cap\mathcal{U}_{j}}^{\top}\mathbf{X}_{\mathcal{U}_{i}\cap\mathcal{U}_{j}}.$
@@
$\|\mathbf{W}^{(i)\top}\mathbf{W}^{(i,j)}\mathbf{W}^{(j)}-\mathbf{I}\|\leq\frac{2\|\mathbf{F}\|}{\sigma_{\min}(\mathbf{X}_{\mathcal{U}_{i}\cap\mathcal{U}_{j}}^{\top}\mathbf{X}_{\mathcal{U}_{i}\cap\mathcal{U}_{j}})}.$ (D.28)
Indeed, we suppose $\mathbf{X}_{\mathcal{U}_{i}\cap\mathcal{U}_{j}}^{\top}\mathbf{X}_{\mathcal{U}_{i}\cap\mathcal{U}_{j}}$ is invertible in Theorem 1. Now suppose $\|\mathbf{F}\|<\sigma_{\min}(\mathbf{X}_{\mathcal{U}_{i}\cap\mathcal{U}_{j}}^{\top}\mathbf{X}_{\mathcal{U}_{i}\cap\mathcal{U}_{j}})$. Then $(\hat{\mathbf{X}}^{(i)}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle})^{\top}\hat{\mathbf{X}}^{(j)}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}$ is also invertible and hence, by Theorem 1 in Li (1995) we have
$\begin{split}\|\mathbf{W}^{(i)\top}\mathbf{W}^{(i,j)}\mathbf{W}^{(j)}-\mathbf{I}\|&\leq\frac{2\|\mathbf{F}\|}{\sigma_{\min}((\hat{\mathbf{X}}^{(i)}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle})^{\top}\hat{\mathbf{X}}^{(j)}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle})+\sigma_{\min}(\mathbf{X}_{\mathcal{U}_{i}\cap\mathcal{U}_{j}}^{\top}\mathbf{X}_{\mathcal{U}_{i}\cap\mathcal{U}_{j}})}\leq\frac{2\|\mathbf{F}\|}{\sigma_{\min}(\mathbf{X}_{\mathcal{U}_{i}\cap\mathcal{U}_{j}}^{\top}\mathbf{X}_{\mathcal{U}_{i}\cap\mathcal{U}_{j}})}.\end{split}$
Otherwise if $\|\mathbf{F}\|\geq\sigma_{\min}(\mathbf{X}_{\mathcal{U}_{i}\cap\mathcal{U}_{j}}^{\top}\mathbf{X}_{\mathcal{U}_{i}\cap\mathcal{U}_{j}})$ then, as $\|\mathbf{W}^{(i)\top}\mathbf{W}^{(i,j)}\mathbf{W}^{(j)}-\mathbf{I}\|\leq 2$, Eq. (D.28) holds trivially.

### v4

@@ Appendix E Technical Lemmas
Eq. (E.15) and Eq. (E.16) implies $\|\mathbf{A}^{[h]}-\mathbf{P}\|\lesssim q_{i}^{-1/2}(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})n_{i}^{1/2}$ and $\lambda_{d}(\mathbf{A}^{[h]})\asymp\lambda_{i,\min}$ with high probability. Then by Wedin’s $\sin\Theta$ Theorem (see e.g., Theorem 4.4 in Stewart and Sun (1990)) we have
$\|\sin\Theta(\widehat{\mathbf{U}}^{[h]},\mathbf{U})\|\leq\frac{\|\mathbf{A}^{[h]}-\mathbf{P}\|}{\lambda_{d}(\mathbf{A}^{[h]})-\lambda_{d+1}(\mathbf{P})}\lesssim\frac{(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})n_{i}^{1/2}}{q_{i}^{1/2}\lambda_{i,\min}}$
with high probability. Let $\mathbf{W}^{[h]}$ be the orthogonal Procrustes alignment between $\widehat{\mathbf{U}}^{[h]}$ and $\mathbf{U}$. Then
$\|\widehat{\mathbf{U}}^{[h]\top}\mathbf{U}-\mathbf{W}^{[h]}\|\leq\|\sin\Theta(\widehat{\mathbf{U}}^{[h]},\mathbf{U})\|^{2}\lesssim\frac{(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})^{2}n_{i}}{q_{i}\lambda_{i,\min}^{2}},$ (E.22)
@@
Now let $\mathbf{Z}^{[h]}=\widehat{\mathbf{U}}^{[h]}\widehat{\mathbf{U}}^{[h]\top}\mathbf{U}-\mathbf{U}$. By Eq. (E.22) and Lemma E.5 we obtain
$\|\mathbf{Z}^{[h]}\|_{2\to\infty}$ $\leq\|\widehat{\mathbf{U}}^{[h]}\widehat{\mathbf{U}}^{[h]\top}\mathbf{U}-\widehat{\mathbf{U}}^{[h]}\mathbf{W}^{[h]}\|_{2\to\infty}+\|\widehat{\mathbf{U}}^{[h]}\mathbf{W}^{[h]}-\mathbf{U}\|_{2\to\infty}$
$\leq\|\widehat{\mathbf{U}}^{[h]}\|_{2\to\infty}\cdot\|\widehat{\mathbf{U}}^{[h]\top}\mathbf{U}-\mathbf{W}^{[h]}\|+\|\widehat{\mathbf{U}}^{[h]}\mathbf{W}^{[h]}-\mathbf{U}\|_{2\to\infty}$
@@
with high probability, where the final inequality follows from the fact that, under the conditions in Eq. (D.3) we have $\frac{(\|\mathbf{P}\|_{\max}+\sigma_{i})n_{i}^{1/2}}{q^{1/2}\lambda_{i,\min}\log^{1/2}n_{i}}\lesssim 1$.
Finally, as $\mathbf{e}_{h}$ and $\mathbf{Z}^{[h]}$ are independent, by Lemma E.6 and the assumption $q_{i}n_{i}\gtrsim\log^{2}n_{i}$ we have
@@
Lemma E.9.
Consider the setting in Lemma E.4, we then have
$\|\mathbf{E}^{(i)}(\widehat{\mathbf{U}}^{(i)}\mathbf{W}_{\mathbf{U}}^{(i)}-\mathbf{U}^{(i)})\|_{2\to\infty}\lesssim\frac{d^{1/2}(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})^{2}n_{i}^{1/2}\log n_{i}}{q_{i}\lambda_{i,\min}}$
@@
$\mathbf{e}_{h}^{\top}(\widehat{\mathbf{U}}\mathbf{W}_{\mathbf{U}}-\mathbf{U})$ $=\mathbf{e}_{h}^{\top}(\widehat{\mathbf{U}}\mathbf{W}_{\mathbf{U}}-\mathbf{U})\mathbf{W}_{\mathbf{U}}^{\top}(\mathbf{W}_{\mathbf{U}}-\widehat{\mathbf{U}}^{\top}\mathbf{U})+\mathbf{e}_{h}^{\top}\mathbf{U}\mathbf{W}_{\mathbf{U}}^{\top}(\mathbf{W}_{\mathbf{U}}-\widehat{\mathbf{U}}^{\top}\mathbf{U})$ (E.23)
$+\mathbf{e}_{h}^{\top}(\widehat{\mathbf{U}}\widehat{\mathbf{U}}^{\top}-\widehat{\mathbf{U}}^{[h]}\widehat{\mathbf{U}}^{[h]\top})\mathbf{U}+\mathbf{e}_{h}^{\top}(\widehat{\mathbf{U}}^{[h]}\widehat{\mathbf{U}}^{[h]\top}\mathbf{U}-\mathbf{U}).$
We now bound all terms on the right hand side of Eq. (E.23). For $\mathbf{e}_{h}^{\top}(\widehat{\mathbf{U}}\mathbf{W}_{\mathbf{U}}-\mathbf{U})\mathbf{W}_{\mathbf{U}}^{\top}(\mathbf{W}_{\mathbf{U}}-\widehat{\mathbf{U}}^{\top}\mathbf{U})$, by Lemma E.2 we have
$\|\mathbf{e}_{h}^{\top}(\widehat{\mathbf{U}}\mathbf{W}_{\mathbf{U}}-\mathbf{U})\mathbf{W}_{\mathbf{U}}^{\top}(\mathbf{W}_{\mathbf{U}}-\widehat{\mathbf{U}}^{\top}\mathbf{U})\|$ $\lesssim\|\mathbf{e}_{h}^{\top}(\widehat{\mathbf{U}}\mathbf{W}_{\mathbf{U}}-\mathbf{U})\|\cdot\|\mathbf{W}_{\mathbf{U}}-\widehat{\mathbf{U}}^{\top}\mathbf{U}\|$ (E.24)
$\lesssim\|\mathbf{e}_{h}^{\top}(\widehat{\mathbf{U}}\mathbf{W}_{\mathbf{U}}-\mathbf{U})\|\cdot\frac{(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})^{2}n_{i}}{q_{i}\lambda_{i,\min}^{2}}$
$=O(\|\mathbf{e}_{h}^{\top}(\widehat{\mathbf{U}}\mathbf{W}_{\mathbf{U}}-\mathbf{U})\|)$
with high probability as $\frac{(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})n_{i}^{1/2}}{q_{i}^{1/2}\lambda_{i,\min}}\lesssim 1$ as implied by Eq. (D.3).
For $\mathbf{e}_{h}^{\top}\mathbf{U}\mathbf{W}_{\mathbf{U}}^{\top}(\mathbf{W}_{\mathbf{U}}-\widehat{\mathbf{U}}^{\top}\mathbf{U})$, by Lemma E.1 and Lemma E.2 we have
$\|\mathbf{e}_{h}^{\top}\mathbf{U}\mathbf{W}_{\mathbf{U}}^{\top}(\mathbf{W}_{\mathbf{U}}-\widehat{\mathbf{U}}^{\top}\mathbf{U})\|$ $\leq\|\mathbf{E}\mathbf{U}\|_{2\to\infty}\cdot\|\mathbf{W}_{\mathbf{U}}-\widehat{\mathbf{U}}^{\top}\mathbf{U}\|$ (E.25)
@@
For $\mathbf{e}_{h}^{\top}(\widehat{\mathbf{U}}\widehat{\mathbf{U}}^{\top}-\widehat{\mathbf{U}}^{[h]}\widehat{\mathbf{U}}^{[h]\top})\mathbf{U}$, by Lemma E.1 and Lemma E.7 we have
[…]
@@
For $\mathbf{e}_{h}^{\top}(\widehat{\mathbf{U}}^{[h]}\widehat{\mathbf{U}}^{[h]\top}\mathbf{U}-\mathbf{U})$, by Lemma E.8 we have
$\|\mathbf{e}_{h}^{\top}(\widehat{\mathbf{U}}^{[h]}\widehat{\mathbf{U}}^{[h]\top}\mathbf{U}-\mathbf{U})\|\lesssim\frac{d^{1/2}(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})^{2}n_{i}^{1/2}\log n_{i}}{q_{i}\lambda_{i,\min}}$ (E.27)
@@
Combining Eq. (E.23), Eq. (E.24), …, Eq. (E.27) we finally obtain
$\|\mathbf{e}_{h}^{\top}(\widehat{\mathbf{U}}\mathbf{W}_{\mathbf{U}}-\mathbf{U})\|$ $\lesssim\frac{d^{1/2}(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})^{2}n_{i}^{1/2}\log n_{i}}{q_{i}\lambda_{i,\min}}$
with high probability as $\frac{(\|\mathbf{P}\|_{\max}+\sigma_{i})n_{i}^{1/2}}{q_{i}^{1/2}\lambda_{i,\min}\log^{1/2}n_{i}}\lesssim 1$ as implied by Eq. (D.3). ∎
E.5 Proof of Lemma D.2
@@
$\mathbf{W}^{(i)\top}\mathbf{W}^{(i,j)}\mathbf{W}^{(j)}=\underset{\mathbf{O}\in\mathcal{O}_{d}}{\arg\min}\|\widehat{\mathbf{X}}^{(i)}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}\mathbf{W}^{(i)}\mathbf{O}-\widehat{\mathbf{X}}^{(j)}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}\mathbf{W}^{(j)}\|_{F}.$
@@
$\mathbf{F}:=\mathbf{W}^{(i)\top}(\widehat{\mathbf{X}}^{(i)}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle})^{\top}\widehat{\mathbf{X}}^{(j)}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}\mathbf{W}^{(j)}-\mathbf{X}_{\mathcal{U}_{i}\cap\mathcal{U}_{j}}^{\top}\mathbf{X}_{\mathcal{U}_{i}\cap\mathcal{U}_{j}}.$
@@
$\|\mathbf{W}^{(i)\top}\mathbf{W}^{(i,j)}\mathbf{W}^{(j)}-\mathbf{I}\|\leq\frac{2\|\mathbf{F}\|}{\sigma_{\min}(\mathbf{X}_{\mathcal{U}_{i}\cap\mathcal{U}_{j}}^{\top}\mathbf{X}_{\mathcal{U}_{i}\cap\mathcal{U}_{j}})}.$ (E.28)
Indeed, we suppose $\mathbf{X}_{\mathcal{U}_{i}\cap\mathcal{U}_{j}}^{\top}\mathbf{X}_{\mathcal{U}_{i}\cap\mathcal{U}_{j}}$ is invertible in Theorem 1. Now suppose $\|\mathbf{F}\|<\sigma_{\min}(\mathbf{X}_{\mathcal{U}_{i}\cap\mathcal{U}_{j}}^{\top}\mathbf{X}_{\mathcal{U}_{i}\cap\mathcal{U}_{j}})$. Then $(\widehat{\mathbf{X}}^{(i)}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle})^{\top}\widehat{\mathbf{X}}^{(j)}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}$ is also invertible and hence, by Theorem 1 in Li (1995) we have
$\begin{split}\|\mathbf{W}^{(i)\top}\mathbf{W}^{(i,j)}\mathbf{W}^{(j)}-\mathbf{I}\|&\leq\frac{2\|\mathbf{F}\|}{\sigma_{\min}((\widehat{\mathbf{X}}^{(i)}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle})^{\top}\widehat{\mathbf{X}}^{(j)}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle})+\sigma_{\min}(\mathbf{X}_{\mathcal{U}_{i}\cap\mathcal{U}_{j}}^{\top}\mathbf{X}_{\mathcal{U}_{i}\cap\mathcal{U}_{j}})}\leq\frac{2\|\mathbf{F}\|}{\sigma_{\min}(\mathbf{X}_{\mathcal{U}_{i}\cap\mathcal{U}_{j}}^{\top}\mathbf{X}_{\mathcal{U}_{i}\cap\mathcal{U}_{j}})}.\end{split}$
Otherwise if $\|\mathbf{F}\|\geq\sigma_{\min}(\mathbf{X}_{\mathcal{U}_{i}\cap\mathcal{U}_{j}}^{\top}\mathbf{X}_{\mathcal{U}_{i}\cap\mathcal{U}_{j}})$ then, as $\|\mathbf{W}^{(i)\top}\mathbf{W}^{(i,j)}\mathbf{W}^{(j)}-\mathbf{I}\|\leq 2$, Eq. (E.28) holds trivially.